OpenStudy (anonymous):

Simplify the expression. (3x^3/4) (8x^1/2) A) 24x^5/4 B) 24x^3/8 C) 11x^2/3 D) 11x^1/4

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

i think B

OpenStudy (jdoe0001):

base is the same, different exponents keep the base sum the exponents

OpenStudy (anonymous):

so its B? 4x2 and 3x1.

OpenStudy (jdoe0001):

\(\Large \bf \left(3x^{\frac{3}{4}}\right)\left(8x^{\frac{1}{2}}\right)\implies 3\cdot 8\cdot x^{\frac{3}{4}+\frac{1}{2}}\)

OpenStudy (anonymous):

but if u add the fractions it wouldnt be a answer

OpenStudy (jdoe0001):

woudln't be an answer? youi mean none of the choices would match?

OpenStudy (anonymous):

same

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

so it has to be B

OpenStudy (jdoe0001):

well.. what did you get from \(\bf \large \left(3x^{\frac{3}{4}}\right)\left(8x^{\frac{1}{2}}\right)\implies 3\cdot 8\cdot x^{\frac{3}{4}+\frac{1}{2}} ?\)

OpenStudy (ageta):

it is B

OpenStudy (ageta):

3x^3(8x^1) =3x3*8x =3*x^3*8*x =(3*8)*(x^3*x) =24*x^4 =24x^4

OpenStudy (ageta):

than you multiply 4*2 it is 8 so it is B☻☻☻☻

OpenStudy (anonymous):

thnx

OpenStudy (ageta):

your welcome

OpenStudy (anonymous):

@jdoe you get 24x^2/3

OpenStudy (anonymous):

http://usatestprep.com/modules/gallery/files/96/9641/9641.png Which triangles are similar? A) ΔABE ∼ ΔACD B) ΔABE ∼ ΔADC C) ΔABE ∼ ΔCBE D) ΔABE ∼ ΔECD

OpenStudy (anonymous):

it's either Aor B

OpenStudy (ageta):

i mean it is A

OpenStudy (ageta):

ya it is A